OpenStudy (zachnovo15):

MEDAL AND FAN. NEED HELP WITH THIS!!!! A biologist is comparing the growth of a population of flies per week to the number of flies a lizard will consume per week. She has devised an equation to solve for which day (x) the lizard would be able to eat the entire population. The equation is 3x = 5x − 1. Explain to the biologist how she can solve this on a graph using a system of equations.

OpenStudy (alekos):

the equation is written incorrectly

OpenStudy (zachnovo15):

3^2=5x-1 @alekos

OpenStudy (alekos):

still not right

OpenStudy (zachnovo15):

It's not 2. It's 3^x=5x-1 @alekos

OpenStudy (zachnovo15):

This is what it shows.

OpenStudy (alekos):

thats better!

OpenStudy (zachnovo15):

Can you help? I'm confused on how to explain it.

OpenStudy (alekos):

yep. The growth of flies is governed by the equation y=3^x and the consumption of flies is y=5x-1. So start by graphing these two equations.

OpenStudy (alekos):

The day that all the flies are consumed is the point of intersection of these two equations

OpenStudy (zachnovo15):

So my graphs look like this

OpenStudy (alekos):

yes, good job using desmos. so the first point of intersection looks like x = 0.5 (half a day). Just click on that point and it'll give you the x value

OpenStudy (zachnovo15):

Ok. So whats next?

OpenStudy (alekos):

thats it, the solution is the 1st point of intersection of the two equations

OpenStudy (zachnovo15):

Ok. Thank you so much!!!!

OpenStudy (alekos):

no problem you're welcome